How to find the amplitude of the wave with equation y=asin (wt-kx) ?

Grade:11

1 Answers

Arun
6 years ago
As we know
Sin²x = (1- cos 2x)/2
Hence when we change this equation of wave
We get
y = A/2 (1- cos 2(wt-kx))
y = A/2 - A/2 sin (pi/2 -(2wt-2kx))
Hence 
Amplitude will be 
A/2